# Sentiment Reasoning for Healthcare

### v3

@@ 6 Conclusion
In this work, we introduce a new task - Sentiment Reasoning - for both speech and text modalities, along with the framework and dataset. In Sentiment Reasoning, given an input transcript (either human transcript or ASR transcript), the model learns to output the sentiment label (POSITIVE, NEUTRAL, or NEGATIVE) and its rationale (the reason for this label). It comprises of two tasks: Sentiment Classification and Rationale Generation.
We thoroughly evaluate the use of rationale during training to improve our models’ interpretability and performance. We found that rationale-augmented training improves model performance in Sentiment Classification in both human and ASR transcripts (1% increase in both accuracy and macro-F1). We also found that the generated rationales generally have different vocabulary to that of human but with similar semantics. Finally, we found no significant difference in the semantic quality of generated rationales between human and ASR transcripts.
@@ 7 Limitations
This study utilized the hybrid ASR system, which is generally recognized as superior in performance compared to the attention-based encoder-decoder or end-to-end ASR systems Lüscher et al. (2019); Prabhavalkar et al. (2023); Raissi et al. (2023). However, the hybrid ASR requires multiple steps, beginning with acoustic feature extraction and progressing through GMM-HMM modeling before transitioning to DNN-HMM modeling, which complicates reproducibility for non-experts.
Additionally, we employed a cascaded speech Sentiment Reasoning approach. This approach uses a previously trained ASR model to generate transcripts that are subsequently input into a LM for downstream sentiment classification and reasoning tasks. Consequently, the weights in the ASR model remain unchanged while the LM weights are updated. In spoken language processing, cascaded approach is prefered due to its straightforwardness, simplicity and superior accuracy Lu (2023); Bentivogli et al. (2021); Tran et al. (2022b); Tseng et al. (2023). Future work should consider the end-to-end sentiment analysis task, where weights in both the ASR model and LM are updated simultaneously, as it holds promise for improved performance.
8 Acknowledgement
We would like to thank Anh Totti Nguyen at Auburn University and Jerry Ngo at MIT for giving insightful feedback.
@@ A.2 ASR-based Speech Sentiment Analysis
Speech sentiment analysis on ASR transcripts is a field that aims to interpret and classify sentiments conveyed in spoken language. As technology advances, ASR systems have become increasingly proficient at transcribing spoken words into text with high accuracy Schneider et al. (2019); Baevski et al. (2020, 2019); Wang et al. (2021b); Chen et al. (2022); Wang et al. (2021a), providing a rich source of data for sentiment analysis. Sentiment analysis algorithms then analyze the transcribed text from speech signal, utilizing language models as decoders to detect positive, negative, or neutral sentiments Lu et al. (2020); Shon et al. (2021); Wu et al. (2022); Tashev and Emmanouilidou (2019); Kaushik et al. (2017).
@@ B.3 Annotation Guidelines
The definition of ”sentiment” encompasses both ”emotions” and ”facts” in our work. Existing works Chen et al. (2020) use both emotions and facts for sentiment labeling. To ensure consistency, our TESOL-certificated professional linguist has developed an initial guideline inspired by Chen et al. (2020) and revised it frequently if necessary as followed:
@@
B.4 Data Quality Control
During the independent annotation process conducted by three annotators, we observed a low inter-annotator agreement, a common occurrence in real-world datasets as noted by Chen et al. (2020). To address this issue, we implemented an alternative label merging approach. We convened a discussion meeting involving the three annotators and two reviewers (one professional linguist and one with a biomedical background). Each annotator was required to justify their chosen sentiment label and its corresponding rationale. A label and its rationale were selected based on the consensus of all three annotators and two reviewers, rather than a majority vote, as employed in other studies Aziz and Dimililer (2020); Saleena et al. (2018).
B.5 Data Statistics Discussion
@@
The NEUTRAL category is the most predominant, accounting for a significant portion of the dataset. With 3802 instances for both train and test set, NEUTRAL sentiments make up approximately half of the dataset. This prevalence of NEUTRAL sentiment is expected, as also seen by a real-world conversational dataset Chen et al. (2020), given the nature of medical consultations, which often involve objective descriptions, explanations, and advice. The NEGATIVE category is the second most common, with around 2395 instances. NEGATIVE sentiments include discussions about serious diseases, negative emotions, and adverse medical outcomes. The substantial presence of negative sentiments reflects the medical context, where discussions about illnesses and symptoms are common. The POSITIVE category, while the least common, still represents a significant portion of the dataset with 1681 instances. POSITIVE sentiments typically involve discussions about recovery processes, positive outcomes, and favorable emotions.
B.6 Data Samples
Table 5 shows 9 examples with 3 samples per sentiment label in our dataset.
@@ Appendix C Details about Experimental Setups
C.1 Training Setup
@@ Appendix D Results on English subset
We use Google Translate to translate our data into English. To ensure that our translated data is usable, we randomly sampled 50 transcripts and check their quality. We further train English models on this English subset of our dataset to ensure full usability.
@@
Appendix E Error Analysis

### v4

@@ 6 Conclusion
In this work, we introduce a new task - Sentiment Reasoning  - for both speech and text modalities, along with the framework and the world’s largest multimodal sentiment analysis dataset. In Sentiment Reasoning, given an input transcript (human transcript or ASR transcript), the model learns to output the sentiment label (POSITIVE, NEUTRAL, or NEGATIVE) and its rationale (the reason for this label). It comprises of two tasks: Sentiment Classification and Rationale Generation.
We menticulously evaluate the use of rationale during training to improve our models’ interpretability and performance. We found that rationale-augmented training improves model performance in Sentiment Classification in both human and ASR transcripts ( +2% increase in both accuracy and macro-F1). We found that the generated rationales have different vocabulary to human rationale but with similar semantics. Finally, we found no major difference in the semantic quality of generated rationales between human and ASR transcripts.
7 Acknowledgement
We thank Anh Totti Nguyen at Auburn University and Jerry Ngo at MIT for insightful feedback.
@@ 8 Limitations
Hybrid ASR: This study utilized the hybrid ASR system, which is generally recognized as superior in performance compared to the attention-based encoder-decoder or end-to-end ASR systems Lüscher et al. (2019); Prabhavalkar et al. (2023); Raissi et al. (2023). However, the hybrid ASR requires multiple steps, beginning with acoustic feature extraction and progressing through GMM-HMM modeling before transitioning to DNN-HMM modeling, which complicates reproducibility for non-experts.
Cascaded speech sentiment analysis approach: While we do report the results for end-to-end systems, our main focus in this paper is on cascaded speech sentiment analysis for Sentiment Reasoning  . This approach uses a previously trained ASR model to generate ASR transcripts that are subsequently input into a language model (LM) for downstream Sentiment Classification and Rationale Generation tasks. Consequently, the weights in the ASR model remain unchanged while the LM weights are updated. In this setting, only semantic features from speech are utilized, omitting other trainable acoustic features, like prosody, tones, etc. In spoken language processing, where semantic features play a more important role than other acoustic features, cascaded approach is prefered due to its straightforwardness, simplicity and superior accuracy Lu (2023); Bentivogli et al. (2021); Tran et al. (2022b); Tseng et al. (2023). Future work should consider the end-to-end sentiment analysis task, where weights in both the ASR model and LM are updated simultaneously, as it might hold promise for improved performance.
@@ A.2 ASR-based Speech Sentiment Analysis
Speech sentiment analysis on ASR transcripts is a field that aims to interpret and classify sentiments conveyed in spoken language. As technology advances, ASR systems have become increasingly proficient at transcribing spoken words into text with high accuracy Schneider et al. (2019); Baevski et al. (2020, 2019); Wang et al. (2021b); Chen et al. (2022); Wang et al. (2021a), providing a rich source of data for sentiment analysis. Sentiment analysis algorithms then analyze the transcribed text from speech signal, utilizing language models as decoders to detect positive, negative, or neutral sentiments Lu et al. (2020); Shon et al. (2021a); Wu et al. (2022); Tashev and Emmanouilidou (2019); Kaushik et al. (2017).
@@ B.3 Annotation Guidelines
The definition of ”sentiment” encompasses both ”emotions” and ”facts” in our work. Existing works Chen et al. (2020); Mohammad (2016); Shon et al. (2021b, 2022, 2023) use both emotions and facts for sentiment labeling.
  • Emotion: Existing literature includes “emotion” as part of “sentiment” Chen et al. (2020); Shon et al. (2021b); Mohammad (2016) and sentiment analysis can be considered a more abstract level of emotion recognition, e.g. polarity of emotions Mohammad (2016).
  • Facts: Many sentiment analysis systems require statements that describe events/situations to be given a sentiment label Chen et al. (2020); Mohammad (2016).
The annotation task consists of two primary steps. First, annotators are required to perform Sentiment Classification. Second, annotators are instructed to provide a rationale behind each class (Rationale Generation ).
To ensure consistency, our TESOL-certificated professional linguist has developed an initial guideline inspired by Chen et al. (2020), which was also adopted by various well-known works Shon et al. (2022, 2023), and revised it frequently if necessary as followed:
B.3.1 Output Annotation
@@
B.4 Annotation Flowchart
Inspired by the well-known annotation flowchart provided by Chen et al. (2020), we asked annotators to adopt the annotation flowchart and we ,if necessary, revised as follows:
  1. 1.
    Does the segment exhibit distinct emotional cues indicative of sentiment, such as laughter for positive affect or yelling for negative affect?
    • Yes – Annotate the corresponding class and also note that:
      • (a) In some instances, individuals may laugh to mitigate the discomfort associated with delivering negative statements. In such cases, it should be classified as neutral.
      • (b) If individuals exhibit a sneer (a smile or laughter with a mocking tone), the corresponding sentiment should be classified as negative in such instances.
    • No - Jump into Step 2
  2. 2.
    Does the segment provide an objective account of the facts?
    • Yes - If the segment lists several positive attributes (e.g., good progress in medical treatment, good signs of health improvement), it is classified as positive. Conversely, if it lists several negative attributes, it is classified as negative. In the absence of a clear preponderance of either, the segment is considered neutral.
    • No - Jump into Step 3
  3. 3.
    Does the segment exhibit a preference?
    • Yes - If the subjective opinion or preference conveys a like or dislike, or expresses a positive (e.g., ”it is beneficial that…”) or negative sentiment, it should be annotated accordingly.
    • No - It’s neutral
  4. 4.
    If the utterance is insufficient in length to accurately assess sentiment, it should be classified as neutral.
B.5 Data Imbalance Discussion
@@
As shown in Table 1, NEUTRAL category is the most predominant, accounting for a significant portion of the dataset. With 3802 instances for both train and test set, NEUTRAL sentiments make up approximately half of the dataset. This prevalence of NEUTRAL sentiment is expected, as also seen by a real-world conversational dataset Chen et al. (2020), given the nature of medical consultations, which often involve objective descriptions, explanations, and advice. The NEGATIVE category is the second most common, with around 2395 instances. NEGATIVE sentiments include discussions about serious diseases, negative emotions, and adverse medical outcomes. The substantial presence of negative sentiments reflects the medical context, where discussions about illnesses and symptoms are common. The POSITIVE category, while the least common, still represents a significant portion of the dataset with 1681 instances. POSITIVE sentiments typically involve discussions about recovery processes, positive outcomes, and favorable emotions.
A slight bias in the distribution of the labels towards NEUTRAL in our dataset (49.94% in the train set, 43.88% in the test set) reflects the nature of real-world medical conversations, rather than a weakness of our work. For context, in comparable real-world sentiment analysis datasets such as Switchboard-Sentiment Chen et al. (2020), the distribution is as follows: 30.4% of the speech segments are labelled as POSITIVE, 17% of the segments are labelled as NEGATIVE, and 52.6% of the segments are labelled as NEUTRAL.
To address this labeling bias issue, future works can leverage techniques for fine-tuning models in data imbalance regimes, such as focal loss Ross and Dollár (2017), class weighting King and Zeng (2001).
B.6 Translation Annotation Process and Translation Quality Control
The data were initially translated from the source language into target languages (many-to-many) using the Gemini Large Language Model (LLM). Following the annotation process by zheng2023judging, the LLM-generated translated transcripts were treated as outputs from a real human annotator. In the data quality process, five human annotators manually corrected and then cross-verified all these translations based on the context of the whole conversation. Only transcripts that received consensus approval from multiple annotators were retained, resulting in an inter-annotator agreement of 100%.
All human annotators possessed a professional language proficiency of C1 or higher (or HSK5 for Chinese) in their respective working languages. Additionally, each annotator had completed basic medical training and demonstrated substantial knowledge of medical terminology in their selected language. Furthermore, they were either currently pursuing or had completed undergraduate or graduate studies in countries where their chosen language is predominantly spoken.
B.7 Data Samples
Table 7 shows 9 examples with 3 samples per sentiment label in our dataset. As the Vietnamese transcripts are obtained from short-formed audio, the transcripts contain characteristics of spoken language which serve as noises to the model (e.g. stuttering, hesitation, etc). In our English translation, we aim to retain these properties, leading to unnatural, incomplete sentence with broken wording.
Figure 3 shows 3 examples per sentiment label for all languages: Vietnamese, English, Chinese (Simplified and Traditional), German and French.
@@ B.7 Data Samples
Figure 3: Some samples from our dataset with versions all available languages.
@@ Appendix C Details about Experimental Setups
C.1 Details of ASR Experiments
C.2 Training Setup
@@ Appendix D Results on English subset
We randomly sampled 50 transcripts and check their quality. We further train English models on this English subset of our dataset to ensure full usability.
@@
Appendix E Results on end-to-end audio language models
E.1 Encoder-Based
PhoWhisper Le et al. (2024): Whisper Radford et al. (2022) trained on an 844-hour Vietnamese dataset. It achieves the state-of-the-art performance on various Vietnamese ASR datasets. To perform Sentiment Classification, we attach a classification head to the encoder layer of the model. We use the PhoWhisper base version in our experiments.
E.2 Audio LLMs
Qwen2-Audio Chu et al. (2024): the state-of-the-art Qwen large audio-language model with 7B parameters, pretrained on diverse audio datasets. We use the Instruct version of this model. We finetune Qwen2-Audio on both Label Only and Label + Rationale settings in our experiments. It should be note that the rationales are text-only.
Appendix F Error Analysis